slope intercept form with the points 4,2 and is perpendicular to 4x+3y=-6
Answer by mananth(16946)

4 x + 3 y = 6
Find the slope of this line

3 y = -4 x + 6
Divide by 3
y = -1.333 x + 2
Compare this equation with y=mx+b
slope m = -1.33

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be 0.75

m= 3/4 ,point ( 4 , 2 )
Find b by plugging the values of m & the point in
y=mx+b
2 = 3.00 + b
b= -1
m= 3/4
The required equation is y = 3/ 4 x-1
m.ananth@hotmail.ca